OpenStudy (anonymous):

A wooden artifact from an ancient tomb contains 45 percent of the carbon-14 that is present in living trees. How long ago, to the nearest year, was the artifact made? (The half-life of carbon-14 is 5730 years.)

OpenStudy (anonymous):

The whole half life every \(k\) years, is expressed with: \[ \left(\frac 12\right)^{t/k} \]

OpenStudy (anonymous):

Because then when \(t=k\), you get: \[ \left(\frac{1}{2}\right)^{k/k} = \frac 12 \]

OpenStudy (anonymous):

Your amount \(a\) at time \(t\), can be expressed as\[ a(t) = a_0\left(\frac{1}{2}\right)^{t/k} \]Where \(a_0=a(0)\), the initial amount.

OpenStudy (anonymous):

In our case, we want it so that \(a(t) = 45\%\cdot a_0\)

OpenStudy (anonymous):

So, putting it all together: \[ 0.45a_0 = a_0\left(\frac{1}{2}\right)^{t/5730} \]And we just solve for \(t\).

OpenStudy (anonymous):

i don't know how to solve for t

OpenStudy (anonymous):

Do you know what a logarithm is?

OpenStudy (anonymous):

this problem is due in 5 min…i just need to know the answer and then ill figure out how i got it

OpenStudy (anonymous):

Oh hell no.

OpenStudy (anonymous):

Is it homework or a test?

OpenStudy (anonymous):

hw

OpenStudy (anonymous):

Taking logarithm of both sides gives: \[ \log_2(0.45) = -\frac{t}{5730} \implies t = -5370\log_2(0.45) \]

OpenStudy (anonymous):

so the answer is?

OpenStudy (jhannybean):

Solve it yourself, please.

OpenStudy (dan815):

(1/2)^(t/5730) = 45 t=?? this also gives same answer da/dt=-ka, suppose u are losing some factor k ln a= -kt +c A=ce^(-kt) A(0)=100 A(5730)=50 A=100*e^(-ln 2/5730 * t ) 45=100*e^(-ln 2/5730 * t ) t=?

OpenStudy (anonymous):

I actually gave the answer, I just didn't approximate it. LOL

OpenStudy (anonymous):

Though my answer doesn't quite make sense... let me double check it.

OpenStudy (anonymous):

Nevermind, I think my answer makes sense now.

OpenStudy (dan815):

ye

OpenStudy (dan815):

e^(-ln 2/5730 * t ) = (e^-ln2)^(t/5730) = (1/2)^(t/5730)